Daily amusement park headcount:

Your boss wishes to know whether or not there is a preferred day of the week for attendance among summer visitors. 700 summer month customers were randomly selected and the day of their visit was recorded. Data are below:

Sunday: 102 Monday: 78 Tuesday: 93 Wednesday: 84 Thursday: 97 Friday: 110 Saturday: 136

a) Conduct a hypothesis test to see if one day of the week is preferred over another.

b) Include the hypothesis, test statistic, level of significance, decision, and conclusion.
Answer #1
observed frequencey, O expected proportion expected frequency,E (O-E)²/E
102 0.143 100.00 0.040
78 0.143 100.00 4.840
93 0.143 100.00 0.490
84 0.143 100.00 2.560
97 0.143 100.00 0.090
110 0.143 100.00 1.000
136 0.143 100 12.960

Ho: not any day of the week is preferred over another.

Ha :one day of the week is preferred over another.

chi square test statistic,X² = Σ(O-E)²/E =   21.980              
                  
level of significance, α=   0.05              
Degree of freedom=k-1=   7   -   1   =   6
                  
P value =   0.0012   [ excel function: =chisq.dist.rt(test-stat,df) ]          
Decision: P value < α, Reject Ho  

one day is preffered over another day

............................

Please revert back in case of any doubt.

Please upvote. Thanks in advance.